# Massive Higher-Spin Fields in the Fractional Quantum Hall Effect

### v1

@@ Abstract
This field can be decomposed into the different helicity modes as $\Phi_{\mu\nu\rho\sigma}=\{\beta\,,\alpha_{a}\,,\phi_{ab}\,,v_{abc}\,,S_{abcd}\}$, where each component is completely symmetric and traceless in its spatial indices. The independent components decompose as
@@
with parentheses denoting a symmetrisation, where dividing by the number of terms in the symmetrisation is understood. The decomposition of the remaining components follows from the traceless condition. Except for $\beta$, each component describes two helicity modes. Under boost transformations these components transform, after taking the limit $c\to\infty$, as follows
@@
which shows that we are dealing with a spin-$4$ field. Making the same redefinition (15) as in the spin-$2$ case, we find in the $c\to\infty$ limit the same continuous and discrete symmetries as in the spin-$2$ case except for the spatial rotations. This leads to a planar spin-$4$ Schroedinger action, which has the same form as the planar spin-$2$ Schroedinger action given in Eq. (17) in terms of $\mathcal{Q}_{T}$, which in turn is equivalent to the tetratic LdG Lagrangian (8) when written in complex form.
From massive higher-spins to p-atic LdG actions: The LdG description for p-atic phases in the FQHE can also be recovered from the $c\to\infty$ limit of relativistic actions describing massive higher-spin fields. Let us stress that moving to an off-shell derivation of the Schroedinger actions might provide a rationale for introducing non-linearities and background couplings as limits of their relativistic counterparts (see, e.g., [55, 56] for recent works on massive higher-spin couplings) although complementary on-shell techniques have been explored too [57].
We first show how to recover the spin-$2$ LdG action and then we generalise the analysis to the spin-$4$ case by rescaling the fields as in Eq. (15). Motivated by the description of the spin-$2$ GMP mode in terms of unimodular gravity [25], we choose to work with the action resulting from the coupling of linearised unimodular gravity with auxiliary Stueckelberg fields $B_{\mu}$ and $S$
@@
where $h_{\mu\nu}$ is traceless and $\kappa$ is the gravitational constant, as we are assuming that this action results from the linearisation of a massive gravity theory. This strategy to build quadratic actions for massive fields is the analogue of that described, e.g., in [58, 59] and we provide more details in the Supplemental Material (SM). The action (22) is invariant under the gauge transformations
@@
As discussed in the SM, the equations of motion following from this action reduce to the FP one (10). We can also gauge away the scalar field $S$ using its algebraic gauge transformation. This leads to a transverse condition $\partial\cdot\epsilon=0$ for the vector parameter, implying that only residual volume-preserving diffeomorphisms in $2+1$ dimensions are allowed. We now decompose $h_{\mu\nu}$ as in (11) in terms of the spatial tensors $S_{ab}$, $v_{a}$ and $\phi$ and we introduce $B_{\mu}=(v,S_{a})$. We also introduce the complex fields $\mathcal{S}_{[2]}=\mathcal{S}_{N}=S_{11}+iS_{12}$, $\mathcal{S}_{[1]}=S_{1}+iS_{2}$ and $\mathcal{V}_{[1]}=v_{1}+iv_{2}$. We then define
@@
where $k$ denotes the spin of the field and $\mathcal{Q}_{[2]}$ has to be identified with the $\mathcal{Q}_{N}$ introduced in Eq. (1). Integrating out the field $\mathcal{U}_{[1]}$ and taking the limit $c\to\infty$ leads to
@@
where we omitted all terms that do not depend on $\mathcal{Q}_{[2]}$ and thus decouple (see SM). In the non-relativistic limit we thus recover the nematic LdG Lagriangian (4) with parameters: $k_{1}=-8\kappa m$, $k_{2}=4\kappa$, $k_{3}=2\kappa mE_{0}$.
Now we turn our attention to a relativistic massive spin-$4$ field  [^2] A non-relativistic limit similar to that discussed below was considered in [83] employing, however, a different action and a different rescaling of the fields.. The analogue of the Stueckelberg Lagrangian (48) involves the Maxwell-like kinetic terms that have been introduced in [61, 62] to generalize linearised unimodular gravity and reads
@@
where $\mu(k)$ stands for a set of $k$ symmetrised indices, while repeated indices denote a symmetrisation. This action is invariant under
$\delta\varphi_{\mu(k)}=k\,\partial_{\mu}\xi_{\mu(k-1)}-\frac{2}{2k-1}\binom{k}{2}\eta_{\mu\mu}\partial\cdot\xi_{\mu(k-2)}+a_{k}\,\xi_{\mu(k)},$ (28)
where $\xi_{\mu\nu\rho}$ and $\xi_{\mu\nu}$ are traceless tensors and the gauge parameters satisfy the constraints
$\partial\cdot\xi_{\mu(k-1)}+b_{k-1}\,\xi_{\mu(k-1)}=0,$ (29)
@@
with fixed $a_{k}$ and $b_{k}$ (see SM). We then decompose the rank-$4$ field $\varphi_{\mu(4)}\equiv\Phi_{\mu\nu\rho\sigma}$ in traceless spatial components as in (19) and we introduce the complex variables $\mathcal{S}_{[4]}=\mathcal{S}_{T}=S_{1111}+iS_{1112}$ and $\mathcal{V}_{[3]}=v_{111}+iv_{112}$. We also introduce a similar decomposition for the auxiliary fields. After redefining the fields as in (25) and integrating out from the action the field $\mathcal{U}_{[3]}$, the field $\mathcal{Q}_{[4]}$ (which coincides with $\mathcal{Q}_{T}$ in Eq. (6)) decouples from the other fields in the limit $c\rightarrow\infty$ (see SM). Its dynamics is described by the Lagrangian in Eq. (8) with the identifications $k^{\prime}_{1}=-8\kappa m$, $k^{\prime}_{2}=4\kappa$, $k^{\prime}_{3}=2\kappa mE_{0}$.
@@
Conclusions and outlook: In this work, we have shown and discussed the relation between p-atic phases and non-relativistic higher-spin fields in the FQHE. We have derived the effective field theories for the higher-spin collective modes in the Laughlin states, given by planar spin-p Schroedinger actions, from $(2+1)$-dimensional relativistic massive higher-spin theories. There are several important open questions that we aim to investigate in future work: i) the effective-field-theory description of higher-spin collective modes in the presence of a curved background [25, 63, 64, 65, 66]. For instance, it would be interesting to see whether one can construct a planar spin-p Schroedinger action in a Newton-Cartan background, generalising the spin-00 analysis of [67]; ii) a supergeometric extension to describe the spin-$3/2$ massive mode [68, 69, 70] and possible further semi-integer higher-spin modes; iii) higher-spin interactions allowing to describe bulk modes away from the isotropic point. For all these developments, one could also exploit the $c\to\infty$ limit of the plethora of relativistic higher-spin interactions that only exist in three dimensions. These have been mainly studied for massless higher-spin fields (see [71] for a review), but investigations of massive interactions appeared, e.g., in [72, 56]; iv) more general higher-spin field theories possibly describing incompressibility and collective modes in higher-dimensional FQH fluids [73]. Finally, we wish to mention that in three dimensions there exist alternative descriptions of the dynamics of massive helicities in terms of higher-derivative theories, for both the spin-$2$ and higher-spin cases; see, e.g., [74, 75, 76, 72, 57, 77] and references therein. We defer to future work an analysis of our non-relativistic limit within this approach.
@@
Acknowledgements: A.C. is a research associate of the Fonds de la Recherche Scientifique – FNRS. His work was supported by FNRS under Grants F.4503.20 and T.0047.24. P.S.-R. was supported by the Austrian Science Fund (FWF), projects P 33789 and P 36619, and by the Norwegian Financial Mechanism 2014-2021 via the Narodowe Centrum Nauki (NCN) POLS grant 2020/37/K/ST3/03390. We thank the Erwin Schrödinger International Institute for Mathematics and Physics and P.S.-R. thanks the University of Mons for hospitality. We are pleased to acknowledge discussions with D. Francia, Z. Papić and S. Pu.
Supplemental Material
Appendix A Proca-like actions for massive particles of any spin
@@
We presented quadratic actions describing relativistic massive spin-$s$ fields that involve a number of auxiliary fields. The rationale for building these actions is that a massive spin-$s$ field in $D$ spacetime dimensions propagates the same helicities as a tower of massless fields of spin comprised between 00 and $s$, with each spin appearing once. As a result, one can build a quadratic action for a massive spin-$s$ field by coupling the free actions for massless fields with spin comprised between 00 and $s$ while preserving a deformation of their free gauge symmetry.
@@
${\cal L}=-\frac{1}{4}\,F^{\mu\nu}F_{\mu\nu}-\frac{m^{2}c^{2}}{2}\,B_{\mu}B^{\mu},\qquad F_{\mu\nu}=\partial_{\mu}B_{\nu}-\partial_{\nu}B_{\mu},$ (30)
does not have any gauge symmetry, but one can reinstate the gauge symmetry of the Maxwell Lagrangian by introducing a “Stueckelberg field” via the shift
@@
and it is manifestly invariant under the gauge transformations
$\delta B_{\mu}$ $=\partial_{\mu}\lambda\,,$ (33a)
$\delta S$ $=mc\,\lambda\,.$ (33b)
The gauge transformation of the scalar field $S$ is algebraic: this means that one can use it to set the field to zero and go back to the action (30). We denote all fields with an algebraic gauge transformation as Stueckelberg fields. We obtained the Lagrangian in Eq. (32) performing a shift in the Proca Lagrangian. Alternatively, one can also consider the most general coupling of a Maxwell field with a free scalar and fix the coefficients demanding that the action be invariant under the transformations (33). This strategy was used in Ref. [58] to describe a massive spin-$s$ field via the coupling of $s$ Fronsdal’s Lagrangians [84], describing massless fields of spin $0\leq k\leq s$. The same outcome can be obtained by recovering the massive theory from the Kaluza-Klein reduction of a spin-$s$ Fronsdal’s action in one more dimension, although in this case one might need to consider field redefinitions in order to diagonalize the resulting kinetic terms; see, e.g., [85]. Fixing completely the algebraic Stueckelberg symmetries then gives the Singh-Hagen Lagrangian [86], which describes a massive spin-$s$ field in terms of a traceless rank-$s$ tensor and a number of auxiliary fields growing with $s$.
We used a strategy similar to that employed in [58], but substituting Fronsdal’s Lagrangians with Maxwell-like ones [61, 62]. The $s=2$ instance of the latter corresponds to the linearization of the Lagrangian of unimodular gravity and, in general, they provide an alternative description of the free dynamics of massless particles in terms of traceless fields. Besides being motivated by the role played by unimodular gravity in the FQHE [25], our choice has also the advantage of allowing one to work with traceless fields while still preserving a gauge symmetry. Working with traceless fields is convenient for defining the non-relativistic limits along the lines of [53], while the gauge symmetry (that is absent in the Singh-Hagen approach) might be useful to introduce systematically non-linearities; see, e.g., [55, 56].
Within this framework, the Lagrangian describing a massive spin-$s$ field contains $s$ traceless fields that we distinguish by the number of indices they carry: $\{\varphi_{\mu(k)}\}$ with $0\leq k\leq s$. We recall that the shorthand $\mu(k)$ denotes a set of $k$ symmetrised indices; for instance, $\varphi_{\mu(2)}:=\varphi_{\mu_{1}\mu_{2}}$. The Lagrangian reads
${\cal L}=\kappa\sum_{k=0}^{s}\left\{\frac{1}{2}\,\varphi^{\mu(k)}\left(\Box\varphi_{\mu(k)}-k\,\partial_{\mu}\partial\cdot\varphi_{\mu(k-1)}-m_{k}\,(mc)^{2}\,\varphi_{\mu(k)}\right)+c_{k}\,mc\,\varphi^{\mu(k)}\partial_{\mu}\varphi_{\mu(k-1)}\right\},$ (34)
@@
$m_{k}=-\frac{(s-k-1)(D+s+k-2)}{D+2k-2},\qquad c_{k}=\sqrt{\frac{k(s-k+1)(D+s+k-4)}{D+2k-4}}\,.$ (35)
Here and in the following, repeated covariant or contravariant indices denote a symmetrization, where dividing by the number of terms in the symmetrization is understood. For instance, $\partial_{\mu}\xi_{\mu}:=\frac{1}{2}\left(\partial_{\mu_{1}}\xi_{\mu_{2}}+\partial_{\mu_{2}}\xi_{\mu_{1}}\right)$. The action corresponding to (34) is invariant under the gauge transformations
$\delta\varphi_{\mu(k)}=k\,\partial_{\mu}\xi_{\mu(k-1)}-\frac{2}{D+2k-4}\binom{k}{2}\eta_{\mu\mu}\partial\cdot\xi_{\mu(k-2)}+c_{k+1}\,mc\,\xi_{\mu(k)}\,,$ (36)
where the gauge parameters $\{\xi_{\mu(k)}\}$ with $10\leq k\leq s-$ 1 are traceless and satisfy the constraints
$\partial\cdot\xi_{\mu(k-1)}-\frac{(D+2k-2)c_{k+1}}{(k+1)(D+2k-4)}\,mc\,\xi_{\mu(k-1)}=0\,,\qquad 1\leq k\leq s-1\,.$ (37)
The Lagrangian (34) depends on a single free mass parameter $m$ and its coefficients can be fixed by demanding invariance under gauge transformations of the form (36). In the limit $m\to 0$ one obtains the sum of $s$ Maxwell-like Lagrangians, which are invariant under gauge transformations generated by traceless and divergenceless gauge parameters, thus presenting themselves as the natural higher-spin generalization of unimodular gravity (linearised volume-preserving diffeomorphisms are indeed generated by transverse vectors).
@@
$\Box h_{\mu\nu}-2\,\partial_{(\mu}\partial\cdot h_{\nu)}-\frac{2}{D}\,\eta_{\mu\nu}\partial\cdot\partial\cdot h-m^{2}c^{2}h_{\mu\nu}+c_{2}\,mc\left(\partial_{(\mu}B_{\nu)}-\frac{1}{D}\,\eta_{\mu\nu}\partial\cdot B\right)=0\,,$ (38)
$\Box B_{\mu}-\partial_{\mu}\partial\cdot B-c_{2}\,mc\,\partial\cdot h_{\mu}=0\,,$ (39)
where we renamed the fields to distinguish them more easily. Notice that the equations of motion must be traceless as the fields entering the Lagrangian. For instance, the double divergence of $h_{\mu\nu}$ appears in Eq. (38) as a result of the traceless projection. Taking a divergence of Eq. (39) one obtains $\partial\cdot\partial\cdot h=0$ and substituting this result in the divergence of Eq. (38) one obtains

### v2

@@ B Tetratic order parameter
This field can be decomposed into the different helicity modes as $\Phi_{\mu\nu\rho\sigma}=\{\beta\,,\alpha_{a}\,,\phi_{ab}\,,v_{abc}\,,S_{abcd}\}$, where each component is completely symmetric and traceless in its spatial indices, so that it describes two helicity modes except for $\beta$. The independent components of $\Phi_{\mu\nu\rho\sigma}$ decompose in analogy with (12) as
@@
with parentheses denoting a symmetrization, where dividing by the number of terms in the symmetrization is understood. The decomposition of the remaining components follows from the traceless condition $\eta_{\mu\nu}\Phi^{\mu\nu\rho\sigma}=0$. Under boost transformations these components transform, after taking the limit $c\to\infty$, as follows
@@
which shows that we are dealing with a spin-$4$ field. Making the same redefinition (16) as in the spin-$2$ case,
$\mathcal{S}_{T}=e^{-i(mc^{2}-E_{0})t}\mathcal{Q}_{T},$ (23)
we find in the $c\to\infty$ limit the same continuous and discrete symmetries as in the spin-$2$ case except for the spatial rotations. Looking for an action invariant under these symmetries leads to a planar spin-$4$ Schroedinger action for $\mathcal{Q}_{T}$, which has the same form as the planar spin-$2$ Schroedinger action given in Eq. (18). The latter in turn is equivalent to the tetratic LdG Lagrangian (8) written in complex form. A similar analysis applies to the spin-$s$ FP equations  [^2] See also [91] and references therein for a group-theoretical discussion of the non-relativistic limit of higher-spin representations of the three-dimensional Poincaré group..
IV From massive higher-spin actions to p-atic LdG actions
The LdG description for p-atic phases in the FQHE can also be recovered from the $c\to\infty$ limit of relativistic actions describing massive higher-spin fields. Let us stress that, aside from providing a complementary viewpoint on our previous derivation, moving to an off-shell derivation of the Schroedinger actions might provide a rationale for introducing non-linearities and background couplings as limits of their relativistic counterparts (see, e.g., [57, 58] for recent works on massive higher-spin couplings), although complementary on-shell techniques have been explored too [59]. As in the previous section, we first show how to recover the spin-$2$ LdG action and then we generalize the analysis to the spin-$4$ case.
A Nematic order parameter
We wish to rederive Eq. (18) starting from the non-relativistic limit of an action leading to the FP equations of motion (11). Motivated by the description of the spin-$2$ GMP mode in terms of unimodular gravity [25], we choose to work with the action resulting from the coupling of linearized unimodular gravity with auxiliary Stueckelberg fields $B_{\mu}$ and $S$,
@@ A Nematic order parameter
where $h_{\mu\nu}$ is traceless and $\kappa$ is the gravitational constant, as we are assuming that this action results from the linearization of a massive gravity theory. This strategy to build quadratic actions for massive fields is the analogue of that described, e.g., in [60, 61] and we provide more details in App. B. This action is invariant under the gauge transformations
@@
As discussed in App. B, the equations of motion derived from the action (24) reduce to the FP ones, i.e., to (11). We can also gauge away the scalar field $S$ using its algebraic gauge transformation. This leads to a transverse condition $\partial\cdot\epsilon=0$ for the vector parameter, implying that only residual volume-preserving diffeomorphisms in $2+1$ dimensions are allowed.
We now decompose $h_{\mu\nu}$ as in (12) in terms of the spatial tensors $S_{ab}$, $v_{a}$ and $\phi$ and we introduce $B_{\mu}=(v,S_{a})$. We also introduce the complex fields
$\begin{split}\mathcal{S}_{[2]}&=\mathcal{S}_{N}=S_{11}+iS_{12},\quad\mathcal{S}_{[1]}=S_{1}+iS_{2},\\ \mathcal{V}_{[1]}&=v_{1}+iv_{2}.\end{split}$ (27)
We then define
@@
where $k$ denotes the rank of the field, while $m$ coincides with the mass parameter in the action (24) and $E_{0}$ is an arbitrary constant. The resulting complex field $\mathcal{Q}_{[2]}$ has to be identified with the $\mathcal{Q}_{N}$ introduced in Eq. (1). Integrating out the field $\mathcal{U}_{[1]}$ and taking the limit $c\to\infty$ leads to
@@
where we omitted all terms that do not depend on $\mathcal{Q}_{[2]}$ and thus decouple (see App. C). In the non-relativistic limit we thus recover the nematic LdG Lagriangian (4) with parameters: $k_{1}=-8\kappa m$, $k_{2}=4\kappa$, $k_{3}=2\kappa mE_{0}$.
B Tetratic order parameter
We now turn our attention to a relativistic massive spin-$4$ field  [^3] A non-relativistic limit similar to that discussed below was considered in [92] employing, however, a different action and a different rescaling of the fields.. The analogue of the Stueckelberg Lagrangian (24) involves the Maxwell-like kinetic terms that have been introduced in [63, 64] to generalize linearized unimodular gravity and reads
@@ B Tetratic order parameter
where $\kappa$ is the gravitational constant, $\mu(k)$ stands for a set of $k$ symmetrized indices, while repeated indices denote a symmetrization. The coefficients in Eq. (30) are fixed by demanding the invariance of the action under a deformation of the gauge symmetry of each kinetic term (see App. B). We then decompose the rank-$4$ field $\varphi_{\mu(4)}\equiv\Phi_{\mu\nu\rho\sigma}$ in traceless spatial components as in (20) and we introduce the complex variables $\mathcal{S}_{[4]}=\mathcal{S}_{T}=S_{1111}+iS_{1112}$ and $\mathcal{V}_{[3]}=v_{111}+iv_{112}$. We also introduce a similar decomposition for the auxiliary fields of rank $k<4$. After redefining the fields as in Eq. (28) and integrating out from the action the field $\mathcal{U}_{[3]}$, the field $\mathcal{Q}_{[4]}$ (which coincides with $\mathcal{Q}_{T}$ in Eq. (6)) decouples from the other fields in the limit $c\rightarrow\infty$ (see App. B for more details). Its dynamics is then described by the Schroedinger Lagrangian in Eq. (8), with the identifications $k^{\prime}_{1}=-8\kappa m$, $k^{\prime}_{2}=4\kappa$, $k^{\prime}_{3}=2\kappa mE_{0}$.
V Measurement proposal for the spin-4 mode
Recently, an inelastic light scattering experiment using circularly polarized photons has measured the chirality of the GMP mode in Gallium Arsenide (GaAs) quantum wells [27], confirming its geometric interpretation as a massive graviton-like mode. Here, we propose a generalization of the experimental approach in [27], to detect the spin-4 mode. According to angular momentum conservation, the angular momentum transferred to a FQH liquid equals the change in photon spin during light scattering. For instance, a GMP mode with spin $s=-2$ is excited when the incident photon has spin $-1$ and the scattered photon has spin $+1$. Formally, to detect a collective higher-spin mode with spin $s=-4$, by adopting a similar approach, we would need to employ an incident polarized graviton with spin $-2$ and measure a scattered polarized graviton with spin $+2$. Clearly, there are no current experiments involving real gravitons. However, effective gravitons could be emulated with light by following the method presented in [65], in which it has been theoretically shown that a circularly polarized gravitational wave can be emulated by two circularly polarized electromagnetic waves with parallel momenta. In other words, two circularly polarized photons with the same helicity and parallel momenta will be required in an inelastic light scattering experiment to detect the spin-4 mode in FQH states  [^4] We wish to thank Lingjie Du for inspiring discussions about this proposal.
VI Conclusions and outlook
@@
We have shown and discussed the relation between p-atic phases and non-relativistic higher-spin fields in the FQHE. We have derived the effective field theories for the higher-spin collective modes in the Laughlin states, given by planar spin-p Schroedinger actions, from the non-relativistic limit of $(2+1)$-dimensional relativistic massive higher-spin theories. There are several important open questions that we aim to investigate in future work: i) the effective-field-theory description of higher-spin collective modes in the presence of a curved background [25, 67, 68, 69, 70]. Indeed, by coupling the higher-spin fields to a background geometry, we expect to obtain novel geometric terms beyond the Wen-Zee term. Once the higher-spin fields are integrated out, the resulting effective action will lead, in addition to the Wen-Zee term induced by the spin-2 mode [14, 71], to extra terms similar to those derived in Ref. [36]. Since these new terms uniquely depend on the probing fields and the background geometry, they can, in principle, identify measurable quantities. Moreover, it would be interesting to see whether one can construct a planar spin-p Schroedinger action in a Newton-Cartan background, generalizing the spin-00 analysis of [72]; ii) a supergeometric extension to describe the spin-$3/2$ massive mode [73, 74, 75] and possible further semi-integer higher-spin modes; iii) higher-spin interactions allowing to describe bulk modes away from the isotropic point. For all these developments, one could also exploit the $c\to\infty$ limit of the plethora of relativistic higher-spin interactions that only exist in three dimensions. These have been mainly studied for massless higher-spin fields (see [76] for a review), but investigations of massive interactions appeared, e.g., in [77, 58]; iv) more general higher-spin field theories possibly describing incompressibility and collective modes in higher-dimensional FQH fluids [78]. Finally, we wish to mention that in three dimensions there exist alternative descriptions of the dynamics of massive helicities in terms of higher-derivative theories, for both the spin-$2$ and higher-spin cases; see, e.g., [79, 80, 81, 77, 59, 82] and references therein. We defer to future work an analysis of our non-relativistic limit within this approach.
@@
Acknowledgements: A.C. is a research associate of the Fonds de la Recherche Scientifique – FNRS. His work was supported by FNRS under Grants F.4503.20 and T.0047.24. P.S.-R. was supported by the Austrian Science Fund (FWF), projects P 33789 and P 36619, and by the Norwegian Financial Mechanism 2014-2021 via the Narodowe Centrum Nauki (NCN) POLS grant 2020/37/K/ST3/03390. We thank the Erwin Schrödinger International Institute for Mathematics and Physics and P.S.-R. thanks the University of Mons for hospitality. We are pleased to acknowledge discussions with D. Francia, Z. Papić and S. Pu, L. Du and B. Yang.
Appendix A Transformations of the relativistic fields under Lorentz boosts
We recall here the Poincaré transformations of the relativistic fields of spin two and four on which we built our symmetry driven derivation of the Schroedinger actions in Eqs. (4) and (8).
We begin from the spin-$2$ field $h_{\mu\nu}$, which transforms under translations and Lorentz transformations with infinitesimal parameters $a^{\mu}$ and $\Lambda_{\mu\nu}=-\Lambda_{\nu\mu}$ as
$h_{\mu\nu}^{\prime}(x^{\prime})=\frac{\partial x^{\rho}}{\partial x^{\prime\mu}}\frac{\partial x^{\sigma}}{\partial x^{\prime\nu}}\,h_{\rho\sigma}(x)$ (31)
with
[…]
This leads to the infinitesimal variations
$\delta h_{\mu\nu}(x)=h^{\prime}_{\mu\nu}(x)-h_{\mu\nu}(x)=\xi^{\lambda}\partial_{\lambda}h_{\mu\nu}+2\partial_{(\mu}\xi^{\rho}h_{\nu)\rho},$ (33)
where $(\mu_{1}\cdots\mu_{s})$ denotes a normalized symmetrization of the enclosed indices, e.g., $A_{(\mu}B_{\nu)}:=\frac{1}{2}\left(A_{\mu}B_{\nu}+A_{\nu}B_{\mu}\right)$. In terms of the spatial traceless components introduced in Eq. (12), these transformations read
$\delta\phi$ $=\xi^{\lambda}\partial_{\lambda}\phi-2\Lambda^{a}v_{a},$ (34a)
$\delta v_{a}$ $=\xi^{\lambda}\partial_{\lambda}v_{a}+\Lambda_{a}{}^{b}v_{b}-\Lambda^{b}S_{ab}+O(c^{-2}),$ (34b)
$\delta S_{ab}$ $=\xi^{\lambda}\partial_{\lambda}S_{ab}+2\Lambda_{(a}{}^{c}S_{b)c}+O(c^{-2}),$ (34c)
where $a,b\in\{1,2\}$ and we denoted the generator of Lorentz boosts as $\Lambda_{a}:=\Lambda_{a0}$, while $\Lambda_{ab}$ is the generator of spatial rotations. These variations show that in the limit $c\to\infty$ the chain of transformations in Eq. (13) holds. In particular, this implies that the symmetric traceless field $S_{ab}$ forms by itself a representation of the Galilei group, which is obtained as a $c\to\infty$ contraction of the Poincaré group as mentioned in the main text.
A spin-$4$ field transforms instead under translations and Lorentz transformations generated by the same infinitesimal parameters as
$h_{\mu\nu\rho\sigma}^{\prime}(x^{\prime})=\frac{\partial x^{\alpha}}{\partial x^{\prime\mu}}\frac{\partial x^{\beta}}{\partial x^{\prime\nu}}\frac{\partial x^{\gamma}}{\partial x^{\prime\rho}}\frac{\partial x^{\delta}}{\partial x^{\prime\sigma}}\,h_{\mu\nu\rho\sigma}(x),$ (35)
with $x^{\prime\mu}$ given again by Eq. (32). This implies the infinitesimal variations
$\delta h_{\mu\nu\rho\sigma}=\xi^{\lambda}\partial_{\lambda}h_{\mu\nu\rho\sigma}+4\partial_{(\mu}\xi^{\lambda}h_{\nu\rho\sigma)\lambda},$ (36)
that in terms of the traceless components introduced in Eq. (20) read
$\delta\beta$ $=\xi^{\lambda}\partial_{\lambda}\beta-4\Lambda^{a}\alpha_{a},$ (37)
$\delta\alpha_{a}$ $=\xi^{\lambda}\partial_{\lambda}\alpha_{a}+\Lambda_{a}{}^{b}\alpha_{b}-3\Lambda^{b}\phi_{ab}+O(c^{-2}),$ (38)
$\delta\phi_{ab}$ $=\xi^{\lambda}\partial_{\lambda}\phi_{ab}+2\Lambda_{(a}{}^{c}\phi_{b)c}-2\Lambda^{c}v_{abc}+O(c^{-2}),$ (39)
$\delta v_{abc}$ $=\xi^{\lambda}\partial_{\lambda}v_{abc}+3\Lambda_{(a}{}^{d}\phi_{bc)d}-\Lambda^{d}S_{abcd}+O(c^{-2}),$ (40)
$\delta S_{abcd}$ $=\xi^{\lambda}\partial_{\lambda}S_{abcd}+4\Lambda_{(a}{}^{e}S_{bcd)e}+O(c^{-2}),$ (41)
corresponding to the chain of transformations summarized in Eq. (21).
Appendix B Proca-like actions for massive particles of any spin
@@
In Sect. IV, we presented quadratic actions describing relativistic massive spin-$s$ fields that involve a number of auxiliary fields. The rationale for building these actions is that a massive spin-$s$ field in $D$ spacetime dimensions propagates the same helicities as a tower of massless fields of spin comprised between 00 and $s$, with each spin appearing once. As a result, one can build a quadratic action for a massive spin-$s$ field by coupling the free actions for massless fields with spin comprised between 00 and $s$ while preserving a deformation of their free gauge symmetry.
@@
${\cal L}=-\frac{1}{4}\,F^{\mu\nu}F_{\mu\nu}-\frac{m^{2}c^{2}}{2}\,B_{\mu}B^{\mu},$ (42)
with $F_{\mu\nu}=\partial_{\mu}B_{\nu}-\partial_{\nu}B_{\mu}$, does not have any gauge symmetry. One can however reinstate the gauge symmetry of the Maxwell Lagrangian by introducing a “Stueckelberg field” via the shift
@@
and it is invariant under the gauge transformations
$\delta B_{\mu}=\partial_{\mu}\lambda,\qquad\delta S=mc\,\lambda.$ (45)
The gauge transformation of the scalar field is algebraic: this means that one can use it to eliminate $S$ and go back to the Lagrangian (42). We denote all fields with an algebraic gauge transformation as Stueckelberg fields. We obtained the Lagrangian in Eq. (44) performing a shift in the Proca Lagrangian. Alternatively, one can also consider the most general coupling of a Maxwell field with a free scalar and fix the coefficients demanding that the action be invariant under the transformations (45).
This strategy was used in Ref. [60] to describe a massive spin-$s$ field via the coupling of $s$ Fronsdal’s Lagrangians [83], describing massless fields of spin $0\leq k\leq s$ (see also Ref. [84]). Fixing completely the algebraic Stueckelberg symmetries then gives the Singh-Hagen Lagrangian [85], which describes a massive spin-$s$ field in terms of a traceless rank-$s$ tensor and a number of auxiliary fields growing with $s$.
In this paper, we used a similar strategy, but substituting Fronsdal’s Lagrangians with Maxwell-like ones [63, 64]. Within this framework, the Lagrangian describing a massive spin-$s$ field contains $s$ traceless fields that we distinguish by the number of indices they carry: $\{\varphi_{\mu(k)}\}$ with $0\leq k\leq s$. We recall that the shorthand $\mu(k)$ denotes a set of $k$ symmetrized indices; for instance, $\varphi_{\mu(2)}:=\varphi_{\mu_{1}\mu_{2}}$. The Lagrangian reads
$\begin{split}{\cal L}=\frac{\kappa}{2}\sum_{k=0}^{s}\varphi^{\mu(k)}\Big{(}\Box\varphi_{\mu(k)}-k\,\partial_{\mu}\partial\cdot\varphi_{\mu(k-1)}\\ -m_{k}\,(mc)^{2}\,\varphi_{\mu(k)}+2c_{k}\,mc\,\partial_{\mu}\varphi_{\mu(k-1)}\Big{)},\end{split}$ (46)
@@
$m_{k}$ $=-\frac{(s-k-1)(D+s+k-2)}{D+2k-2},$ (47a)
$c_{k}$ $=\sqrt{\frac{k(s-k+1)(D+s+k-4)}{D+2k-4}}.$ (47b)
Here and in the following, repeated covariant or contravariant indices denote a symmetrization, where dividing by the number of terms in the symmetrization is understood. For instance, $\partial_{\mu}\xi_{\mu}:=\frac{1}{2}\left(\partial_{\mu_{1}}\xi_{\mu_{2}}+\partial_{\mu_{2}}\xi_{\mu_{1}}\right)$. The action involving (46) is invariant under the gauge transformations
$\begin{split}\delta\varphi_{\mu(k)}&=k\,\partial_{\mu}\xi_{\mu(k-1)}+c_{k+1}\,mc\,\xi_{\mu(k)}\\ &-\frac{2}{D+2k-4}\binom{k}{2}\eta_{\mu\mu}\partial\cdot\xi_{\mu(k-2)},\end{split}$ (48)
where the gauge parameters $\{\xi_{\mu(k)}\}$, with $10\leq k\leq s-$ 1, are traceless and satisfy the constraints
$\partial\cdot\xi_{\mu(k-1)}-\frac{(D+2k-2)c_{k+1}}{(k+1)(D+2k-4)}\,mc\,\xi_{\mu(k-1)}=0,$ (49)
with $11\leq k\leq s-$ 1. The Lagrangian (46) depends on a single free mass parameter $m$ and its coefficients can be fixed by demanding invariance under gauge transformations of the form (48). In the limit $m\to 0$ one obtains the sum of $s$ Maxwell-like Lagrangians, which are invariant under gauge transformations generated by traceless and divergenceless gauge parameters, thus presenting themselves as the natural higher-spin generalization of unimodular gravity (linearized volume-preserving diffeomorphisms are indeed generated by transverse vectors).
@@
$\Box h_{\mu\nu}-2\,\partial_{(\mu}\partial\cdot h_{\nu)}-\frac{2}{D}\,\eta_{\mu\nu}\partial\cdot\partial\cdot h-m^{2}c^{2}h_{\mu\nu}$
$\quad+c_{2}\,mc\left(\partial_{(\mu}B_{\nu)}-\frac{1}{D}\,\eta_{\mu\nu}\partial\cdot B\right)=0\,,$ (50a)
$\Box B_{\mu}-\partial_{\mu}\partial\cdot B-c_{2}\,mc\,\partial\cdot h_{\mu}=0\,,$ (50b)
where we renamed the fields as in Eq. (24) to distinguish them more easily. Notice that the equations of motion must be traceless as the fields entering the Lagrangian. For instance, the double divergence of $h_{\mu\nu}$ appears in Eq. (50a) as a result of the traceless projection. Taking a divergence of Eq. (50b) one obtains $\partial\cdot\partial\cdot h=0$ and substituting this result in the divergence of Eq. (50a) one obtains